# SANA-Sprint: One-Step Diffusion with Continuous-Time Consistency Distillation

### v3

@@ Abstract
Abstract: This paper presents SANA-Sprint, an efficient diffusion model for ultra-fast text-to-image (T2I) generation. SANA-Sprint is built on a pre-trained foundation model and augmented with hybrid distillation, dramatically reducing inference steps from 20 to 1-4. We introduce three key innovations: (1) We propose a training-free approach that transforms a pre-trained flow-matching model for continuous-time consistency distillation (sCM), eliminating costly training from scratch and achieving high training efficiency. Our hybrid distillation strategy combines sCM with latent adversarial distillation (LADD): sCM ensures alignment with the teacher model, while LADD enhances single-step generation fidelity. (2) SANA-Sprint is a unified step-adaptive model that achieves high-quality generation in 1-4 steps, eliminating step-specific training and improving efficiency. (3) We integrate ControlNet with SANA-Sprint for real-time interactive image generation, enabling instant visual feedback for user interaction. SANA-Sprint establishes a new Pareto frontier in speed-quality tradeoffs, achieving state-of-the-art performance with 7.59 FID and 0.74 GenEval in only 1 step — outperforming FLUX-schnell (7.94 FID / 0.71 GenEval) while being 10× faster (0.1s vs 1.1s on H100). It also achieves 0.1s (T2I) and 0.25s (ControlNet) latency for 1024$\times$1024 images on H100, and 0.31s (T2I) on an RTX 4090, showcasing its exceptional efficiency and potential for AI-powered consumer applications (AIPC). Code and pre-trained models will be open-sourced.
@@ 1 Introduction
Figure 1: (a) Our SANA-Sprint accelerate the inference speed for generating 1024 $\times$ 1024 images, achieving a remarkable speedup from FULX-Schnell’s 1.94 seconds to only 0.03 seconds. This represents a 64$\times$ improvement over the current state-of-the-art step-distilled model, FLUX-Schnell, as measured with a batch size of 1 on an NVIDIA A100 GPU. The ratio is calculated based on Transformer latency. (b) Additionally, our model demonstrates efficient GPU memory usage during training, outperforming other distillation methods in terms of memory cost. The GPU memory is measured using official code, 1024 $\times$ 1024 images and on a single A100 GPU.
@@
Despite their promise, key limitations hinder broader adoption. GAN-based methods suffer from training instability due to oscillatory adversarial dynamics and mode collapse. GANs face challenges due to the need to map noise to natural images without supervision, making unpaired learning more ill-posed than paired learning, as highlighted in [18, 19]. This instability is compounded by architectural rigidity, which demands meticulous hyperparameter tuning when adapting to new backbones or settings. VSD-based methods involve the joint training of an additional diffusion model, which increases computational overhead and imposes significant pressure on GPU memory, and requires careful tuning [20]. Consistency models, while stable, suffer quality erosion in ultra-few-step regimes (e.g., $<$4 steps), particularly in text-to-image tasks where trajectory truncation errors degrade semantic alignment. These challenges underscore the need for a distillation framework that harmonizes efficiency, flexibility, and quality.
@@
As a result, SANA-Sprint achieves excellent speed/quality tradeoff, benefiting from a hybrid objective, inheriting sCM’s diversity preservation and alignment with the teacher, while integrating LADD’s fidelity enhancement: experiments show a 0.6 lower FID and 0.4 higher CLIP-Score at 2-step generations compared to standalone sCM, with 3.9 lower FID and 0.9 higher CLIP-Score over pure latent adversarial approaches. As shown in Fig. 1, SANA-Sprint achieves state-of-the-art performance in FID and GenEval benchmark, surpassing recent advanced methods including SD3.5-Turbo, SDXL-DMD2, and Flux-schnell. Especially, SANA-Sprint is 64.7$\times$ faster than Flux-Schnell and exceeds in FID (7.59 vs 7.94) and GenEval (0.74 vs 0.71).
Moreover, SANA-Sprint demonstrates unprecedented inference speed—generating 1024$\times$1024 images in 0.31 seconds on a laptop with consumer-grade GPUs (NVIDIA RTX 4090) and 0.1 seconds on H100 GPU, 8.4$\times$ speedup than teacher model SANA. This efficiency unlocks transformative applications that require instant visual feedback: in ControlNet-guided image generation/editing, by integrating with ControlNet, SANA-Sprint enables instant interaction with 250ms latency on H100. SANA-Sprint exhibits robust scalability and is potentially suitable for human-in-the-loop creative workflows, AIPC, and immersive AR/VR interfaces.
@@
  • Excellent Speed/Quality Tradeoff: SANA-Sprint achieves exceptional performance with only 1-4 steps. SANA-Sprint generates a 1024×1024 image in only 0.10s-0.18s on H100, achieving state-of-the-art 7.59 FID on MJHQ-30K and 0.74 GenEval score - surpassing FLUX-schnell (7.94 FID/0.71 GenEval) while being 10$\times$ faster.
@@ 2.1 Diffusion Model and Its Variants
Figure 2: Training paradigm of SANA-Sprint. In SANA-Sprint, we use the student model for synthetic data generation ($\hat{x_{0}}$) and JVP calculation, and we use the teacher model for velocity ($\mathrm{d}x/\mathrm{d}t$) compute and its feature for the GAN loss, which allows us train sCM and GAN together and have only one training model purely in the latent space. Details of training objective and TrigFlow Transformation are in Eq. 9, Eq. 11 and Sec. 3.1.
@@ 2.2 Consistency Models
Although discrete-time CMs work well in practice, the additional discretization errors brought by numerical ODE solvers are inevitable. Continuous-time CMs correspond to the limiting case of $\Delta t\rightarrow 0$ in Eq. 2. When choosing $d(\boldsymbol{x},\boldsymbol{y})=\|\boldsymbol{x}-\boldsymbol{y}\|_{2}^{2}$, the expression simplifies to:
@@ 3.1 Training-Free Transformation to TrigFlow
Score-based generative models (diffusion, flow matching, and TrigFlow) can denoise data with proper data scales and signal-to-noise ratios (SNRs) [^1] For a diffusion model $\boldsymbol{x}_{t}=\alpha_{t}\boldsymbol{x}_{0}+\sigma_{t}\boldsymbol{z}$, SNR is defined as $\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}$ aligned with training. However, flow matching cannot directly denoise TrigFlow-scheduled data due to three mismatches: First, their time domains differ: TrigFlow uses $[0,\frac{\pi}{2}]$, while flow matching is defined on $[0,1]$. Second, their noise schedules are distinct—TrigFlow maintains $\cos^{2}(t_{\texttt{Trig}})+\sin^{2}(t_{\texttt{Trig}})=1$, while flow matching yields $t_{\texttt{FM}}^{2}+(1-t_{\texttt{FM}})^{2}<1$, causing data scale discrepancies. Finally, their prediction targets differ: flow matching predicts $z-x_{0}$ with static coefficients $(1,-1)$, whereas TrigFlow predicts $\cos(t)z-\sin(t)x_{0}$ with time-varying coefficients. These mismatches in temporal parameterization, SNR, and output necessitate explicit input/output transformations.
@@
$\quad+(1-2t_{\texttt{FM}}+2t_{\texttt{FM}}^{2})\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t,\texttt{FM}},t_{\texttt{FM}},\boldsymbol{y})\Big{]}.$
@@
The details and proof of Proposition. 3.1 are in Appendix D. The transformations of both input and output are all differentiable making it compatible with auto differentiation. As validated by Tab. 1, the transformation is lossless in both theory and practice. The training-free transformation is depicted in the gray box of Fig. 2.
Figure 3: Efficient Distillation via QK Normalization, Dense Timestep Embedding, and Training-free Schedule Transformation. (a) We compare gradient norms and visualizations with/without QK Normalization, showing its stabilizing effect. (b) Gradient norm curves for timestep scales (0$\sim$1 vs. 0$\sim$1000) highlight impacts on stability and stability and quality. (c) PCA-based similarity analysis of timestep embeddings. (d) Image results after 5,000 iterations of fine-tuning with (left) and without (right) the proposed schedule transfer (Sec. 3.1).
Table 1: Comparison of original Flow-based SANA model and training-free transformation of TrigFlow-based SANA-Sprint model. We evaluate the FID and CLIP-Score before and after the transformation in Sec. 3.1.
@@
where $\widehat{\boldsymbol{f_{\theta^{-}}}}$ is the parameterized sCM as in Eq. 4 after replacing $\boldsymbol{F_{\theta}}$ with $\widehat{\boldsymbol{F_{\theta}}}$ in Proposition. 3.1, $\boldsymbol{x}_{t},t$ refers to $\boldsymbol{x}_{t,\texttt{Trig}},t_{\texttt{Trig}}$, and $w_{\boldsymbol{\phi}}(t)$ is an adaptive weighting function to minimize variance across different timesteps following [31, 17].
@@ 3.2 Stabilizing Continuous-Time Distillation
Dense Time-Embedding. As analyzed in sCM [17], the instability issues in continuous-time CMs primarily stem from the unstable scale of $\frac{\mathrm{d}\boldsymbol{f_{\theta}}}{\mathrm{d}t}$ in Eq. 9. This instability can be traced back to the expression $\frac{\mathrm{d}F_{\theta^{-}}}{\mathrm{d}t}=\nabla_{\boldsymbol{x}_{t}}F_{\theta^{-}}\frac{\mathrm{d}\boldsymbol{x}_{t}}{\mathrm{d}t}+\partial_{t}F_{\theta^{-}}$ in Eq. 5, which ultimately originates from the time derivative term $\partial_{t}F_{\theta^{-}}$:
@@
In previous flow matching models like SANA [32], SD3 [22], and FLUX [21], the noise coefficient $c_{\text{noise}}(t)=1000t$ amplifies the time derivative $\partial_{t}F_{\theta^{-}}$ by a factor of 1000, leading to significant training fluctuations. To address this, we set $c_{\text{noise}}(t)=t$ and fine-tuned SANA for 5k iterations. As shown in Fig. 3 (b), this adjustment reduce excessively large gradient norms (originally exceeding $10^{3}$) to more stable levels. Furthermore, PCA visualization in Fig. 3 (c) reveals that our dense time-embedding design results in more densely packed and similar embeddings for timesteps between 0$\sim$1. This refinement improved training stability and accelerated convergence over 15k iterations.
QK-Normalization. When scaling up the model from 0.6B to 1.6B, we encounter similar issues with excessively large gradient norms, often exceeding $10^{3}$, which lead to training collapse. To address this, we introduce RMS normalization [33] to the Query and Key in both self- and cross-attention modules of the teacher model during fine-tuning. This modification enhances training stability significantly, even with a brief fine-tuning process of only 5,000 iterations. By using the fine-tuned teacher model to initialize the student model, we achieve a more stable gradient norm, as shown in Fig. 3 (a), thereby making the distillation process viable where it was previously infeasible.
@@ 3.3 Improving Continuous-Time CMs with GAN
The adversarial loss $\mathcal{L}_{\text{adv}}$ is equivalent to the GAN loss shown at the bottom of Fig. 2. In summary, SANA-Sprint combines sCM loss with GAN loss: $\mathcal{L}=\mathcal{L}_{sCM}+\lambda\mathcal{L}_{\text{adv}}$, where $\lambda=0.5$ by default, as in Tab. 6.
@@
In our early experiments, we adopt the timestep sampling distribution of sCM’s generator (student model) for GAN loss, given by $t=\arctan{(\frac{e^{\tau}}{\sigma_{d}})}$, where $\tau\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ with two hyperparameters $P_{\text{mean}}$ and $P_{\text{std}}$. To further enhance one- and few-step generation performance and improve overall generation quality, we introduce an additional weighting at $t=\frac{\pi}{2}$. Specifically, with probability $p$, the training timestep is set to $\frac{\pi}{2}$, while with probability $1-p$, it follows the original timestep sampling distribution of sCM’s generator. We find that this modification significantly improves the model’s capability for one- and few-step generation, as shown in Tab. 6.
@@ 3.4 Application: Real-Time Interactive Generation
For ControlNet tasks, we extract Holistically-Nested Edge Detection (HED) scribbles from input images as conditions to guide image generation. Following PixArt’s [40] design principles, we train our SANA-ControlNet teacher model on 1024$\times$1024 resolution images. During sampling, HED maps serve as additional conditioning inputs to the Transformer model, allowing precise control over image generation while maintaining structural details. Our experiments show that the distilled SANA-Sprint-ControlNet model retains the controllability of the teacher model and achieves fast inference speeds of approximately 200 ms on H100 machines, enabling near-real-time interaction. The effectiveness of our approach is demonstrated in Sec. F.3.
@@ 4.1 Experimental Setup
Our experiments employ a two-phase training strategy, with detailed settings and evaluation protocols outlined in Sec. F.1. The teacher models are pruned and fine-tuned from the larger SANA-1.5 4.8B model [42], followed by distillation using our proposed training paradigm. We evaluate performance using metrics including FID, CLIP Score on the MJHQ-30K [43], and GenEval [44].
@@ 4.2 Efficiency and Performance Comparison
We compare SANA-Sprint with state-of-the-art text-to-image diffusion and timestep distillation methods in Tab. 2 and Fig. 4. Our SANA-Sprint models focus on timestep distillation, achieving high-quality generation with 1-4 inference steps, competing with the 20-step teacher model, as shown in Fig. 5. More details about the timestep setting are given in Sec. F.2.
@@ 4.3 Analysis
Schedule Transfer. To validate the effectiveness of our proposed schedule transfer in Sec. 3.1, we conduct ablation studies on a flow matching model SANA [32], comparing its performance with and without schedule transformation to TrigFlow [17]. As shown in Fig. 3 (d), removing schedule transfer leads to training divergence due to incorrect signals. In contrast, incorporating our schedule transfer enables the model to achieve decent results within 5,000 iterations, demonstrating its crucial role in efficiently adapting flow matching models to TrigFlow-based consistency models.
Influence of CFG Embedding. To clarify the influence of Classifier-Free Guidance (CFG) embedding in our model, we maintain the setting of incorporating CFG into the teacher model, as established in previous works [45, 13, 17]. Specifically, during the training of the student model, we uniformly sample the CFG scale of the teacher model from the set ${4.0,4.5,5.0}$. To integrate CFG embedding [11] into the student model, we add it as an additional condition to the time embedding, multiplying the CFG scale by 0.1 to align with our denser timestep embeddings. We conduct experiments with and without CFG embedding to evaluate its role. As shown in Tab. 6, incorporating CFG embedding significantly improves CLIP score by 0.94.
Effects of sCM and LADD. We evaluate the effectiveness of each component by comparing models trained with only the sCM loss or the LADD loss. As shown in Tab. 6, training with LADD alone results in instability and suboptimal performance, achieving a higher FID score of 12.20 and a lower CLIP score of 27.00. In contrast, combining both sCM and LADD losses improves model performance, yielding a lower FID score of 8.11 and a higher CLIP score of 28.02, demonstrating their complementary benefits. Using sCM alone achieves a FID score of 8.93 and a CLIP score of 27.51, indicating that while sCM is effective, adding LADD further enhances performance. The weighting ablations for sCM and LADD loss are shown in Tab. 6, with additional timestep distribution ablations provided in Sec. F.2
Additional Max-Time Weighting. We validate the proposed max-time weighting strategy in LADD (see Sec. 3.3) through experiments with both sCM and LADD losses. As shown in Tab. 6, this weighting significantly improves performance. We test the strategy at 0%, 50%, and 70% max-time ($t=\pi/2$) probabilities, finding that 50% is the best balance, while higher probabilities provide only marginal gains. However, considering the qualitative results, we finally choose 50% as the default max-time weighting.

### v4

@@ Abstract
Abstract: This paper presents SANA-Sprint, an efficient diffusion model for ultra-fast text-to-image (T2I) generation. SANA-Sprint is built on a pre-trained foundation model and augmented with hybrid distillation, dramatically reducing inference steps from 20 to 1-4. We introduce three key innovations: (1) We propose a training-free approach that transforms a pre-trained flow-matching model for continuous-time consistency distillation (sCM), eliminating costly training from scratch and achieving high training efficiency. Our hybrid distillation strategy combines sCM with latent adversarial distillation (LADD): sCM ensures alignment with the teacher model, while LADD enhances single-step generation fidelity. (2) SANA-Sprint is a unified step-adaptive model that achieves high-quality generation in 1-4 steps, eliminating step-specific training and improving efficiency. (3) We integrate ControlNet with SANA-Sprint for real-time interactive image generation, enabling instant visual feedback for user interaction. SANA-Sprint establishes a new Pareto frontier in speed-quality tradeoffs, achieving state-of-the-art performance with 7.59 FID and 0.74 GenEval in only 1 step — outperforming FLUX-schnell (7.94 FID / 0.71 GenEval) while being 10× faster (0.1s vs 1.1s on H100). It also achieves 0.1s (T2I) and 0.25s (ControlNet) latency for $\times1024$ images on H100, and 0.31s (T2I) on an RTX 4090, showcasing its exceptional efficiency and potential for AI-powered consumer applications (AIPC). Code and pre-trained models will be open-sourced.
@@ 1 Introduction
[…]
@@
Despite their promise, key limitations hinder broader adoption. GAN-based methods suffer from training instability due to oscillatory adversarial dynamics and mode collapse. GANs face challenges due to the need to map noise to natural images without supervision, making unpaired learning more ill-posed than paired learning, as highlighted in [18, 19]. This instability is compounded by architectural rigidity, which demands meticulous hyperparameter tuning when adapting to new backbones or settings. VSD-based methods involve the joint training of an additional diffusion model, which increases computational overhead and imposes significant pressure on GPU memory, and requires careful tuning [20]. Consistency models, while stable, suffer quality erosion in ultra-few-step regimes (e.g., <<4 steps), particularly in text-to-image tasks where trajectory truncation errors degrade semantic alignment. These challenges underscore the need for a distillation framework that harmonizes efficiency, flexibility, and quality.
@@
[…]
@@
[…]
@@ 2.1 Diffusion Model and Its Variants
Figure 2: Training paradigm of SANA-Sprint. In SANA-Sprint, we use the student model for synthetic data generation ($\hat{x_{0}}$) and JVP calculation, and we use the teacher model for velocity ($\mathrm{d}x/\mathrm{d}t$) compute and its feature for the GAN loss, which allows us train sCM and GAN together and have only one training model purely in the latent space. Details of training objective and TrigFlow Transformation are in Eq.˜9, Eq.˜11 and Sec. 3.1.
@@ 2.2 Consistency Models
Although discrete-time CMs work well in practice, the additional discretization errors brought by numerical ODE solvers are inevitable. Continuous-time CMs correspond to the limiting case of $\Delta t\rightarrow 0$ in Eq.˜2. When choosing $d(\boldsymbol{x},\boldsymbol{y})=\|\boldsymbol{x}-\boldsymbol{y}\|_{2}^{2}$, the expression simplifies to:
@@ 3.1 Training-Free Transformation to TrigFlow
Score-based generative models (diffusion, flow matching, and TrigFlow) can denoise data with proper data scales and signal-to-noise ratios (SNRs) [^1] For a diffusion model $\boldsymbol{x}_{t}=\alpha_{t}\boldsymbol{x}_{0}+\sigma_{t}\boldsymbol{z}$, SNR is defined as $\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}$ aligned with training. However, flow matching cannot directly denoise TrigFlow-scheduled data due to three mismatches: First, their time domains differ: TrigFlow uses $0,\frac{\pi}{2}]$, while flow matching is defined on $[0,1]$. Second, their noise schedules are distinct—TrigFlow maintains $\cos^{2}(t_{\texttt{Trig}})+\sin^{2}(t_{\texttt{Trig}})=1$, while flow matching yields $t_{\texttt{FM}}^{2}+(1-t_{\texttt{FM}})^{2}<1$, causing data scale discrepancies. Finally, their prediction targets differ: flow matching predicts $z-x_{0}$ with static coefficients (1,1)(1,-1), whereas TrigFlow predicts $\cos(t)z-\sin(t)x_{0}$ with time-varying coefficients. These mismatches in temporal parameterization, SNR, and output necessitate explicit input/output transformations.
@@
$\qquad+(1-2t_{\texttt{FM}}+2t_{\texttt{FM}}^{2})\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t,\texttt{FM}},t_{\texttt{FM}},\boldsymbol{y})\Big].$
@@
The details and proof of Proposition.˜3.1 are in Appendix˜D. The transformations of both input and output are all differentiable making it compatible with auto differentiation. As validated by Tab.˜1, the transformation is lossless in both theory and practice. The training-free transformation is depicted in the gray box of Fig.˜2.
[…]
Table 1: Comparison of original Flow-based SANA model and training-free transformation of TrigFlow-based SANA-Sprint model. We evaluate the FID and CLIP-Score before and after the transformation in Sec.˜3.1.
@@
where $\widehat{\boldsymbol{f_{\theta^{-}}}}$ is the parameterized sCM as in Eq.˜4 after replacing $\boldsymbol{F_{\theta}}$ with $\widehat{\boldsymbol{F_{\theta}}}$ in Proposition.˜3.1, $\boldsymbol{x}_{t},t$ refers to $\boldsymbol{x}_{t,\texttt{Trig}},t_{\texttt{Trig}}$, and $w_{\boldsymbol{\phi}}(t)$ is an adaptive weighting function to minimize variance across different timesteps following [31, 17].
@@ 3.2 Stabilizing Continuous-Time Distillation
Dense Time-Embedding. As analyzed in sCM [17], the instability issues in continuous-time CMs primarily stem from the unstable scale of $\frac{\mathrm{d}\boldsymbol{f_{\theta}}}{\mathrm{d}t}$ in Eq.˜9. This instability can be traced back to the expression $\frac{\mathrm{d}F_{\theta^{-}}}{\mathrm{d}t}=\nabla_{\boldsymbol{x}_{t}}F_{\theta^{-}}\frac{\mathrm{d}\boldsymbol{x}_{t}}{\mathrm{d}t}+\partial_{t}F_{\theta^{-}}$ in Eq.˜5, which ultimately originates from the time derivative term $\partial_{t}F_{\theta^{-}}$:
@@
[…]
QK-Normalization. When scaling up the model from 0.6B to 1.6B, we encounter similar issues with excessively large gradient norms, often exceeding $10^{3}$, which lead to training collapse. To address this, we introduce RMS normalization [33] to the Query and Key in both self- and cross-attention modules of the teacher model during fine-tuning. This modification enhances training stability significantly, even with a brief fine-tuning process of only 5,000 iterations. By using the fine-tuned teacher model to initialize the student model, we achieve a more stable gradient norm, as shown in Fig.˜3 (a), thereby making the distillation process viable where it was previously infeasible.
@@ 3.3 Improving Continuous-Time CMs with GAN
The adversarial loss $\mathcal{L}_{\text{adv}}$ is equivalent to the GAN loss shown at the bottom of Fig.˜2. In summary, SANA-Sprint combines sCM loss with GAN loss: $\mathcal{L}=\mathcal{L}_{sCM}+\lambda\mathcal{L}_{\text{adv}}$, where $\lambda=0.5$ by default, as in Tab.˜6.
@@
In our early experiments, we adopt the timestep sampling distribution of sCM’s generator (student model) for GAN loss, given by $t=\arctan{(\frac{e^{\tau}}{\sigma_{d}})}$, where $\tau\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ with two hyperparameters $P_{\text{mean}}$ and $P_{\text{std}}$. To further enhance one- and few-step generation performance and improve overall generation quality, we introduce an additional weighting at $t=\frac{\pi}{2}$. Specifically, with probability $p$, the training timestep is set to $\frac{\pi}{2}$, while with probability 1p1-p, it follows the original timestep sampling distribution of sCM’s generator. We find that this modification significantly improves the model’s capability for one- and few-step generation, as shown in Tab.˜6.
@@ 3.4 Application: Real-Time Interactive Generation
For ControlNet tasks, we extract Holistically-Nested Edge Detection (HED) scribbles from input images as conditions to guide image generation. Following PixArt’s [40] design principles, we train our SANA-ControlNet teacher model on $\times1024$ resolution images. During sampling, HED maps serve as additional conditioning inputs to the Transformer model, allowing precise control over image generation while maintaining structural details. Our experiments show that the distilled SANA-Sprint-ControlNet model retains the controllability of the teacher model and achieves fast inference speeds of approximately 200 ms on H100 machines, enabling near-real-time interaction. The effectiveness of our approach is demonstrated in Sec.˜F.3.
@@ 4.1 Experimental Setup
Our experiments employ a two-phase training strategy, with detailed settings and evaluation protocols outlined in Sec.˜F.1. The teacher models are pruned and fine-tuned from the larger SANA-1.5 4.8B model [42], followed by distillation using our proposed training paradigm. We evaluate performance using metrics including FID, CLIP Score on the MJHQ-30K [43], and GenEval [44].
@@ 4.2 Efficiency and Performance Comparison
We compare SANA-Sprint with state-of-the-art text-to-image diffusion and timestep distillation methods in Tab.˜2 and Fig.˜4. Our SANA-Sprint models focus on timestep distillation, achieving high-quality generation with 1-4 inference steps, competing with the 20-step teacher model, as shown in Fig.˜5. More details about the timestep setting are given in Sec.˜F.2.
@@ 4.3 Analysis
Schedule Transfer. To validate the effectiveness of our proposed schedule transfer in Sec.˜3.1, we conduct ablation studies on a flow matching model SANA [32], comparing its performance with and without schedule transformation to TrigFlow [17]. As shown in Fig.˜3 (d), removing schedule transfer leads to training divergence due to incorrect signals. In contrast, incorporating our schedule transfer enables the model to achieve decent results within 5,000 iterations, demonstrating its crucial role in efficiently adapting flow matching models to TrigFlow-based consistency models.
Influence of CFG Embedding. To clarify the influence of Classifier-Free Guidance (CFG) embedding in our model, we maintain the setting of incorporating CFG into the teacher model, as established in previous works [45, 13, 17]. Specifically, during the training of the student model, we uniformly sample the CFG scale of the teacher model from the set 4.0,4.5,5.0{4.0,4.5,5.0}. To integrate CFG embedding [11] into the student model, we add it as an additional condition to the time embedding, multiplying the CFG scale by 0.1 to align with our denser timestep embeddings. We conduct experiments with and without CFG embedding to evaluate its role. As shown in Tab.˜6, incorporating CFG embedding significantly improves CLIP score by 0.94.
Effects of sCM and LADD. We evaluate the effectiveness of each component by comparing models trained with only the sCM loss or the LADD loss. As shown in Tab.˜6, training with LADD alone results in instability and suboptimal performance, achieving a higher FID score of 12.20 and a lower CLIP score of 27.00. In contrast, combining both sCM and LADD losses improves model performance, yielding a lower FID score of 8.11 and a higher CLIP score of 28.02, demonstrating their complementary benefits. Using sCM alone achieves a FID score of 8.93 and a CLIP score of 27.51, indicating that while sCM is effective, adding LADD further enhances performance. The weighting ablations for sCM and LADD loss are shown in Tab.˜6, with additional timestep distribution ablations provided in Sec.˜F.2
Additional Max-Time Weighting. We validate the proposed max-time weighting strategy in LADD (see Sec.˜3.3) through experiments with both sCM and LADD losses. As shown in Tab.˜6, this weighting significantly improves performance. We test the strategy at 0%, 50%, and 70% max-time ($t=\pi/2$) probabilities, finding that 50% is the best balance, while higher probabilities provide only marginal gains. However, considering the qualitative results, we finally choose 50% as the default max-time weighting.